# Alternating Direction Method of Multipliers AR AutoRally

### v1

@@ III-D Performance Recommendations
The performance capabilities of the MPPI-Generic library is highly dependent on the choices of block sizes for the various kernels, Dynamics, and Cost Functions. There is no one choice to be made so instead, we can provide some general rules of thumbs that we have seen work more often than not.
  • Set the thread block $x$ dimension size of both the Dynamics and Cost kernels to be a multiple of 32. This allows all the threads of a warp to be doing the same operation. We have seen in some cases that lowering the thread block $x$ dimension to 16 can provide some improvements but this has not been true for the majority of dynamics and cost functions.
  • Set the thread block $y$ dimension of Dynamics and Cost Functions to the state dim and $1$ respectively. This depends on whether the Dynamics/Cost Function are utilizing the multi-threading capability available to them but for most of the pre-defined Dynamics, they are using the $y$ axis of the thread block to do dynamics updates. Meanwhile, most Cost Functions are not taking advantage of the multi-threading as they are returning a single value, the cost.
  • Keep Dynamics and Cost Function kernels’ thread block sizes low. The limit to the size of a thread block is currently $1024$ so it might be tempting to try to fill that. However, there are various synchronization points in these kernels to ensure that data has been loaded before being used. As the thread block size increases, the more time is spent waiting at these synchronization points in the Dynamics and Cost Function kernels.
  • When developing your own Dynamics, it can be pertinent to make the output dimensions divisible by 2 or 4. This can done by filling the enum with filler values. This allows the code to use the more efficient memory loading GPU instructions discussed in Section III-B and reduce the total number of memory calls by a factor of at least 2. One might be tempted to also ensure that the control dimension is divisible by 4. However, the control dimension is also used to generate samples so while there may be improvements to the efficiency of reading memory, it will also take longer to generate samples that are being read.
  • Minimize the use of if statements in the GPU code when possible. Since GPUs execute a warp at a time, having different threads in the warp following different paths of an if statement can cause warp divergence where every thread ends up executing both code paths by running the if statement twice.
IV API Structure
@@ 34 DYN_T::state_array x = dynamics->getZeroState(); // set up initial state
We will describe the library API and its usage in three stages. At the beginner level, the desire is to use a provided Dynamics, Cost Function, and Controller to control a system. This requires the least amount of code writing on the part of the user as most of the code is already provided. The library user would only need to write a specific Plant class to properly interface with whatever system they are wanting to control and the executable which sets up and runs the controller itself. The intermediate level is where the user might want to implement a dynamics model or cost function that does not exist in the base library. Finally, at the advanced level, we will show how to implement a custom Feedback Controller, Controller, or Sampling Distribution.
@@ 11 SimpleCartpolePlant(std::shared_ptr<CONTROLLER_T> controller, int hz, int optimization_stride)
32 double getCurrentTime() { return current_time_; }
33
34 double getPoseTime() { return this->state_time_; }
35
@@
When using MPPI in \@iaciMPC MPC fashion, we need to use a Plant wrapper around our controller. The Plant houses methods to obtain new data such as state, calculate the optimal control sequence at a given rate using the latest information available, and provide the latest control to the external system while providing the necessary tooling to ensure there are no race conditions. As this is the class that provides the interaction between the algorithm and the actual system, it is a component that has to be modified for every use case. For this example, we will implement a plant inheriting from BasePlant that houses the external system completely internal to the class. Specifically, we will write our plant to run the dynamics inside pubControl() in order to produce a new state. We shall then call updateState() at a different fixed rate from the controller re-planning rate to show that the capability of the code base.
In Lst. 3, we can see a simple implementation of a Plant. SimpleCartpolePlant instantiates a CartpoleDynamics variable upon creation, overwrites the required virtual methods from BasePlant, and sets up the dynamics update to occur within pubControl(). Looking at the constructor, we pass a shared pointer to a Controller, an integer representing the controller replanning rate, and the minimum timestep we want to adjust the control trajectory by when performing multiple optimal control calculations to the base Plant constructor, and then create our stand-in system dynamics. pubControl() is where we send the control to the system and so in this case, we create necessary extra variables and then pass the current state ${\bf x}_{t}$ and control ${\bf u}_{t}$ as prev_state and u respectively to the Dynamics’ step() method to get the next state, ${\bf x}_{t+1}$, in the variable current_state_. We also update the current time to show that system has moved forward in time. Looking at this class, a potential issue arises as it is templated on the controller which in turn might not use CartpoleDynamics as its Dynamics class. This can be easily remedied by replacing any reference to CarpoleDynamics with CONTROLLER_T::TEMPLATED_DYNAMICS to make this plant work with the Dynamics used by the instantiated Controller.
@@
Now that we have written our specialized Plant class, we can then go back to the minimal example in Lst. 2 and make some modifications to use the controller \@iaciMPC MPC fashion. For this example, we will just run a simple for loop that calls the Plant’s runControlIteration() and updateState() methods to simulate a receiving a new state from the system and then calculating a new optimal control sequence from it. The updateState() method calls pubControl() internally so the system state and the current time will get updated as this runs. For real-time scenarios, there is also the runControlLoop() Plant method which can be launched in a separate thread and calls runControlIteration() internally at the specified replanning rate.
@@ IV-B Intermediate
In the previous section, much of the underlying structure of the library was glossed over as it was not necessary to understand. As we get to implementing our own Dynamics or Cost Functions however, there are some basic principles to go over. This library is running code on two different devices, the CPU and the GPU. Some classes, such as Plants and Controllers, don’t have methods that need to run on the GPU whereas other classes like the Dynamics and Cost Functions do. The GPU can do many computations in parallel but in order to properly utilize it, that can require different code than what runs on the CPU. As such, when looking at implementing a new Dynamics or Cost Function, there will be some methods that will have to be implemented twice, once for the CPU and once for the GPU.
@@ IV-B1 Custom Dynamics
Let us first look at Dynamics. The first thing that needs to be implemented for a new Dynamics class is a new parameter structure. We have implemented a dictionary structure in the form of enum to define the state, control, and output vectors and these dictionaries are stored in the parameter. For example, in Lst. 4, we show a basic parameter structure implemented for a unicycle model.
@@ 24 NUM_OUTPUTS
By implementing these enum, we can use these state names later on in the Dynamics and Cost Function to make it clear what state we are referring to. The ending values of NUM_STATES, NUM_CONTROLS, and NUM_OUTPUTS are also used to determine the size of each of the resulting dimensions for creating statically-sized Eigen::Matrix data types such as state_array and control_array.
@@ 42 Eigen::Ref<state_array> x_dot);
Next, we need to implement the basic methods needed to be overwritten. These methods are shown in Lst. 5 and start with getDynamcisModelName() which returns the name of the Dynamics model. Next are the CPU and GPU versions of computeStateDeriv(). The CPU and GPU versions are differentiated firstly by the __device__ keyword at the front of the GPU code to designate that method will only live on the GPU. Next, we also only use Eigen::Matrixf data-types on the CPU and raw float pointers on the GPU. The computeStateDeriv() both would implement the following dynamics,
@@
as shown in Lst. 6. You can see here the use of the enum we created earlier using the S_INDEX() and C_INDEX() macros. This allows us to not need to know the order of the states or controls in the underlying data type and more importantly, these same enum can also be used in Cost classes to ensure that they are compatible with multiple dynamics as long as the dynamics define the same enum.
@@
An important note is that there are separate instances of the class allocated on the CPU and GPU for dynamics and cost functions. In order to update the GPU version you must overwrite paramsToDevice() to copy from the CPU side class to the GPU side class. Parameter structures have a default implementation for this, but if helper classes are added their corresponding copy methods will need to be called.
@@ 15 UnicycleCost(cudaStream_t stream = nullptr);
If we make a Dynamics class, we also generally will want to make a corresponding Cost. While there is a default quadratic cost implementation that allows for use with any dynamics, we will implement a quick Cost class to show the bare minimum to overwrite. For this example, we will choose to think of our unicycle on a road of some width pointing in the $x$ direction. We want to keep the unicycle on the road so we will penalize the absolute $y$ position linearly up to the road width and then penalize it quadratically if it goes outside the road.
To do this, we will first construct a parameter structure for the cost class. It will inherit from the base CostParams class and needs to know the CONTROL_DIM and have a variable for the width of the road and coefficient for the cost. From there, we implement a basic Cost class that has to overwrite computeStateCost() and terminalCost() on both the CPU and GPU sides. It also creates a default constructor and name method getCostFunctionName(). This basic implementation can be seen in Lst. 7. Notice that we have created an alias for our dynamics params struct on line 11. This is to allow us to use the enum defined there inside our cost function as well. Other things to note are that again, our cost is based on the output rather than the state in computeStateCost(). For our simple Unicycle Dynamics, the output is the same as the state.
In Lst. 8, the cost function described previously is implemented. We can use enum macros such O_IND_CLASS() when we are outside of the Dynamics class to still find the appropriate output value. This code can be implemented in both the CPU and GPU computeStateCost() methods and we can choose terminalCost() to return 00.
@@
These new Dynamics and Cost could then be easily incorporated into our previous controller examples and that should be enough to get most people started on using this library. There are more options in the Cost and Dynamics classes to improve GPU performance but those are left to Section III.
@@ IV-C Advanced
In this final section, we will show off how to customize Controllers, Feedback Controllers, and Sampling Distributions. Customizing these allows users to create new sampling-based controllers or change the sampling distributions used in specfic scenarios.
@@ IV-C1 Feedback Controllers
One thing we have not yet discussed is the Feedback Controller API. Feedback Controllers are useful even when using MPPI and become necessary for Tube-MPPI and RMPPI. The only Feedback Controller currently implemented is iLQR but we will show how to construct a simple PID controller for each state/control combination.
@@
Now let us show how to utilize the Feedback Controller API to create a new PID controller. We start by constructing the params structure for the GPU Feedback class, shown in Lst. 9. Note that it is templated off of the Dynamics so that it can create the appropriately-sized arrays for each gain. Since this is a generic PID controller, we will stick to the full feedback matrix representation for each type of gain. Next, we need to implement the GPU Feedback Controller for the PID in Lst. 10. The only method that is required to be overwritten is the k() feedback method but as PIDs require some history to calculate the $i$ and $d$ portions, we will also request extra memory by setting the SHARED_MEM_REQUEST_BLK_BYTES variable. This variable allows for extra shared memory per trajectory sample which is necessary for keeping track of history on the GPU.
@@
28 float* prev_e = theta[(tid * 2 + 1) * DYN_T::STATE_DIM];
@@
33 e_der[i] = (curr_e[i] - prev_e[i]) / this->state_.dt;
@@
After writing the GPU feedback class, we now just have to write the CPU feedback class, shown in Lst. 11. We only have two methods to overwrite here in computeFeedback() and k_(). computeFeedback() is used to calculate new feedback gains given a new trajectory. For our simple PID class, we will stick to persistent PID gains so this can be left empty. The k_() method is just the CPU version of the feedback calculation. For the PID, we can now use member variables to hold the history required to calculate the $i$ and $d$ portions.
@@
68 control_array k_(const Eigen::Ref<const state_array>& x_act,
69 const Eigen::Ref<const state_array>& x_goal,
@@ 78 NUM_SAMPLES * sizeof(float), cudaMemcpyHostToDevice, this->stream_));
List of Listings 13 Basic CEM computeControl() implementation. This method copies the initial state to the GPU, creates the control samples, calculates the state trajectories and costs of each sample, creates the elite set, updates the mean of the sampling distribution and then copies that mean back as the optimal control sequence.
In order to create a new sampling-based Controller, there are a few methods that must be overwritten. We will walk through how to create a Controller using the CEM [5] as our desired algorithm. Succinctly put, the CEM method samples control trajectories from a Gaussian Distribution, and then uses the best $k$ samples, also known as the elite set, to calculate new parameters for the Guassian distribution. For this example, we will simplify it to have constant variance and the elite set is used to update the mean of the distribution. The major CEM parameter is the size of the elite set, so we create a new Params struct in Lst. 12 capturing this as a percentage of the total number of samples. From there, the methods that must be overwritten from the base Controller API are getControllerName(), computeControl(), and calculateSampledStateTrajectories(). getControllerName() just returns the name of the new controller, i.e. CEM. Next, computeControl() is the main method of the Controller class. It takes in the new initial state and calculates the new optimal control sequence from that starting position; an example is shown in Lst. 13. The basic steps are to put the shifted optimal control sequence and initial state onto the GPU, generate the control samples, run the samples through the Dynamics and Cost function, calculate the weights of each sample, use these weights to update the parameters of the sampling distribution, get the new optimal control sequence, and then calculate the corresponding optimal state trajectory. In our example, we add an additional method, calcualteEliteSet(), to find the elite set and zero out the weights of every other sample specifically for CEM. The final method to overwrite, calculateSampledStateTrajectories(), is a method used to generate visualizations based on a subset of sampled trajectories. Users can set a percentage of the sampled trajectories they would like to use for visualization with setPercentageSampledControlTrajectoriesHelper() and this method then can be used to generate state trajectories for those samples and put them on the CPU so that they can then be used for visualization purposes. For this simple example, we have no visualization system to plug this data into so we leave this method empty for our CEM implementation.

### v2

@@ 34 DYN_T::state_array x = dynamics->getZeroState(); // set up initial state
The performance capabilities of the MPPI-Generic library is highly dependent on the choices of block sizes for the various kernels, Dynamics, and Cost Functions. There is no one choice to be made so instead, we provide some general rules of thumbs that we have seen work more often than not.
  • Set the thread block $x$ dimension size of both the Dynamics and Cost kernels to be a multiple of 32. This allows all the threads in a warp to do the same operation. We have seen in some cases that lowering the thread block $x$ dimension to 16 can provide some improvements but this has not been true for the majority of dynamics and cost functions.
  • Set the thread block $y$ dimension for Dynamics equal to the state dimension and Cost Functions to 1. This depends on whether the Dynamics/Cost Function are utilizing the multi-threading capability available to them but for most of the pre-defined Dynamics, they are using the $y$ axis of the thread block to do dynamics updates. Meanwhile, most Cost Functions are not taking advantage of the parallelization as they return a single value, the cost.
  • Keep Dynamics and Cost Function kernels’ thread block sizes low. The limit to the size of a thread block is currently $1024$ so it might be tempting to fill that. However, there are various synchronization points in these kernels to ensure that data has been loaded before being used. As the thread block size increases, more time is spent waiting at these synchronization points in the Dynamics and Cost Function kernels.
  • When developing your own Dynamics, it can be pertinent to make the output dimensions divisible by 2 or 4. This can done by adding extraneous values to the enum. This allows the code to use more efficient memory-loading GPU instructions discussed in Section III-B and reduce the number of memory calls by a factor of at least 2. One might be tempted to also ensure that the control dimension is divisible by 4. However, the control dimension is also used to generate samples so while there may be improvements to the efficiency of reading memory, it will also slow down generating samples.
  • Minimize the occurrence of warp divergence in your GPU code when possible. The most common way these occur is by using if statements based on values that might be different for individual threads in a warp. In some cases, converting the desired bool expression into a float will be faster while still providing the desired branching.
IV API Structure
We describe the library API and its usage in three stages. At the beginner level, the desire is to use a provided Dynamics, Cost Function, and Controller to control a system. This requires the least amount of code writing on the part of the user as most of the code is already provided. The library user would only need to write a specific Plant class to properly interface with whatever system they are wanting to control and the executable which sets up and runs the controller itself. The intermediate level is where the user might want to implement a dynamics model or cost function that does not exist in the base library. Finally, at the advanced level, we show how to implement a custom Feedback Controller, Controller, or Sampling Distribution.
@@ 11 SimpleCartpolePlant(std::shared_ptr<CONTROLLER_T> controller, int hz, int optimization_stride)
29 double getCurrentTime() { return current_time_; }
30 double getPoseTime() { return this->state_time_; }
@@
We start by showing an example of just using a single iteration of MPPI to produce an optimal control sequence in Lst. 2. At the beginning of the example (Lines 8, 9, 10, 11, 12 and 13), we create aliases such DYN_T for CartpoleDynamics to keep the code fairly succinct. The Feedback Controller is not used for this example but it is required to exist. The Sampling Distribution and Controller have parameters that need to be set; their use spans from algorithmic to performance – Lines 22, 23, 26, 27, 28, 29 and 30 respectively. The std_dev, dt_, and lambda_ are parameters affecting the MPPI update rule whereas the dynamics_rollout_dim_ and cost_rollout_dim_ parameters adjust how fast the MPPI algorithm is computed using recommendations from Section III-D. Once the components are initialized, we create an instance of the MPPI Controller, passing the Dynamics, Cost Function, Feedback Controller, Sampling Distribution, and controller parameters to the constructor. On Line 34, we create an initial zero state for the dynamics using getZeroState() and compute an optimal control sequence with computeControl() on Line 35. We return the control sequence as an Eigen::Matrixf with $n_{u}$ rows and $T$ columns using getControlSeq() to print out.
@@
When using MPPI in \@iaciMPC MPC fashion, we need to use a Plant wrapper around our controller. The Plant houses methods to obtain new data such as state, calculate the optimal control sequence at a given rate using the latest information available, and provide the latest control to the external or ground truth system while providing the necessary tooling to ensure there are no race conditions. As this class provides the interaction between the algorithm and the actual system, it is a component that has to be modified for every use case. For Lst. 3, we implement a plant (SimpleCartpolePlant) inheriting from BasePlant that contains the ground truth system completely internal to the class. Specifically, our plant runs the external dynamics inside pubControl() in order to produce a new state. We then call updateState() at a different fixed rate from the controller re-planning rate to show that the capability of the code base. SimpleCartpolePlant instantiates a CartpoleDynamics object in its constructor, overwrites the required virtual methods from BasePlant, and sets up the dynamics update to occur within pubControl(). Looking at the constructor on Line 11, we pass a shared pointer to a Controller, an integer representing the controller replanning rate, and the minimum optimization stride, before creating our stand-in system dynamics. In a new MPC iteration, we shift the start of the mean control sequence we sample around to the maximum between the number of timesteps since the last optimization and the minimum optimization stride. For use in MPC, we recommend setting the minimum optimization stride to $1$. pubControl() on Line 16 is where we send the control to the system. In this case, we create necessary extra variables to pass the current state ${\bf x}_{t}$ and control ${\bf u}_{t}$ as prev_state and u respectively to the Dynamics’ step() method to get the next state, ${\bf x}_{t+1}$, in the variable current_state_. We also update the current time on Line 23 to show the system has moved forward in time. Looking at this class, an issue arises as the Controller it is templated upon might not use CartpoleDynamics as its Dynamics class. This is easily remedied by replacing any reference to CarpoleDynamics with CONTROLLER_T::TEMPLATED_DYNAMICS to make this Plant work with the Dynamics used by the instantiated Controller.
@@
Now that we have written our specialized Plant class, we can make some modifications to Lst. 2 to use the controller in \@iaciMPC MPC fashion. For this example, we would run a simple for loop that calls the Plant’s runControlIteration() and updateState() methods to simulate a receiving a new state from the system and then calculating a new optimal control sequence from it, replacing Line 35. The updateState() method calls pubControl() internally so the system state and the current time would update at each for loop iteration. For real-time scenarios, the runControlLoop() Plant method can be launched in a separate thread and calls runControlIteration() internally at the specified re-planning rate.
@@ IV-B Intermediate
In the previous section, much of the underlying structure of the library was glossed over. As we get to implementing our own Dynamics or Cost Functions however, there are some basic principles to go over. This library is running code on two different devices, the CPU and the GPU. Some classes, such as Plants and Controllers, do not have methods that need to run on both whereas other classes like the Dynamics and Cost Functions do. The GPU can do many computations in parallel but in order to properly utilize it, that can require different code than what runs on the CPU. As such, when looking at implementing a new Dynamics or Cost Function, there are some methods that have to be implemented as two very similar functions, once for the CPU and once for the GPU.
@@ IV-B1 Custom Dynamics
The first thing that needs to be implemented for a new Dynamics class is a new parameter structure. We implemented a dictionary-like structure in the form of enum to define the state, control, and output vectors and these dictionaries are stored in the parameter. For example, in Lst. 4, we show a basic parameter structure implemented for a unicycle model.
@@ 24 NUM_OUTPUTS
By implementing these enum, we can use these state names later on in the Dynamics and Cost Function to make it clear what state we are referring to. The ending values of NUM_STATES, NUM_CONTROLS, and NUM_OUTPUTS are also used to determine the size of each of the resulting dimensions for creating statically-sized Eigen::Matrixf data types such as state_array and control_array.
@@ 42 Eigen::Ref<state_array> x_dot);
Next, we implement the necessary overwritable methods. These methods are shown in Lst. 5 and start with getDynamcisModelName() which returns the name of the Dynamics model. Next are the CPU and GPU versions of computeStateDeriv() on Lines 39 and 44 respectively. The CPU and GPU versions are differentiated firstly by the __device__ keyword at the front of the GPU code to designate that method only runs on the GPU. Next, we also only use Eigen::Matrixf data-types on the CPU and raw float pointers on the GPU. The computeStateDeriv() both would implement the following dynamics,
@@
as shown in Lst. 6. Note the use of the enum we created earlier with the S_INDEX() and C_INDEX() macros. This allows us to not need to know the order of the states or controls in the underlying data type and more importantly, these same enum can also be used in Cost classes to ensure that compatibility with multiple dynamics as long as the dynamics define the necessary enum values.
@@
An important note is that there are separate instances of the class allocated on the CPU and GPU for dynamics and cost functions. In order to update the GPU version, you must overwrite paramsToDevice() to copy from the CPU side class to the GPU side class. Parameter structures have a default implementation for this, but if helper classes are added, their corresponding copy methods will need to be called.
@@ 15 UnicycleCost(cudaStream_t stream = nullptr);
If we make a Dynamics class, we also generally need to make a corresponding Cost. While a default quadratic cost implementation that allows for use with any dynamics exists, we implement a quick Cost class to show what methods to overwrite. For this example, we choose to think of our unicycle on a road of some width pointing in the $x$ direction. We want to keep the unicycle on the road so we penalize the absolute $y$ position linearly up to the road width and penalize it quadratically if it goes outside the road.
To do this, we first construct a parameter structure for the cost class. It inherits from the base CostParams class and needs to know the CONTROL_DIM and have a variable for the width of the road and coefficient for the cost. From there, we implement a basic Cost class that has to overwrite computeStateCost() and terminalCost() on both the CPU and GPU sides. It also creates a default constructor and name method getCostFunctionName(). This basic implementation is shown in Lst. 7. Notice that we created an alias for our dynamics parameter’s struct on Line 13. This is to allow us to use the enum defined there inside our cost function as well. Other things to note are that again, our cost is based on the output rather than the state in computeStateCost(). For our simple Unicycle Dynamics, the output is the same as the state.
@@
In Lst. 8, the cost function described previously is implemented. We can use enum macros such as O_IND_CLASS() when we are outside of the Dynamics class to still find the appropriate output value. This code can be implemented in both the CPU and GPU computeStateCost() methods and we choose terminalCost() to return 00.
These new Dynamics and Cost are easily incorporated into our previous controller examples and that should be enough to get most people started on using this library. There are more options in the Cost and Dynamics classes to improve GPU performance but those are left to Section III.
@@ IV-C Advanced
In this final subsection, we show off how to customize Controllers, Feedback Controllers, and Sampling Distributions. Customizing these allows users to create new sampling-based controllers or change the sampling distributions used in specific scenarios.
@@ IV-C1 Feedback Controllers
Feedback Controllers are useful even when using MPPI and become necessary for Tube-MPPI and RMPPI. The only Feedback Controller currently implemented is iLQR but we show how to construct a simple PID controller for each state/control combination.
@@
Now let us show how to utilize the Feedback Controller API to create a new PID controller. We start by constructing the parameter structure for the GPU Feedback class, shown in Lst. 9. Note that it is templated off of the Dynamics so that it can create the appropriately-sized arrays for each gain. Since this is a generic PID controller, we use the full feedback matrix representation for each type of gain. Next, we need to implement the GPU Feedback Controller for the PID in Lst. 10. The only method that is required to be overwritten is the k() feedback method on Line 23 but as PIDs require some history to calculate the $i$ and $d$ portions, we request extra memory by setting the SHARED_MEM_REQUEST_BLK_BYTES variable on Line 21. This variable allows for extra shared memory per trajectory sample which is necessary for keeping track of history on the GPU.
@@
28 float* pre_e = theta[(tid*2+1) * DYN_T::STATE_DIM];
@@
33 e_der[i] = (curr_e[i] - pre_e[i]) / this->state_.dt;
@@
51template <class DYN_T, int NUM_TIMESTEPS>
52class PIDFeedback : public FeedbackController<gpuPID<DYN_T>, PIDParams, NUM_TIMESTEPS> {
@@
68 control_array k_(const Eigen::Ref<const state_array>& x_act, const Eigen::Ref<const state_array>& x_goal,
@@
After writing the GPU feedback class, we now just have to write the CPU feedback class, shown in Lst. 11. We have two methods to overwrite here in computeFeedback() and k_(). computeFeedback() is used to calculate new feedback gains given a new trajectory. For our simple PID class, we stick to constant PID gains so this can be left empty. The k_() method on Line 68 is just the CPU version of the feedback calculation. For the CPU version of the PID, we can now use member variables to hold the history required to calculate the $i$ and $d$ portions.
@@ 78 NUM_SAMPLES * sizeof(float), cudaMemcpyHostToDevice, this->stream_));
List of Listings 13 Basic CEM computeControl() implementation. This method copies the initial state to the GPU, creates the control samples, calculates the state trajectories and costs of each sample, creates the elite set, updates the mean of the sampling distribution and copies that mean back as the optimal control sequence.
We walk through how to create a new sampling-based Controller using CEM [5] as our desired algorithm. Succinctly put, the CEM method samples control trajectories from a Gaussian Distribution, and uses the best $k$ samples, known as the elite set, to calculate new parameters for the Guassian distribution. For this example, we simplify it to have constant variance and the elite set is used to update the mean of the distribution. The major CEM parameter is the size of the elite set, so we create a new parameter struct in Lst. 12 capturing this as a percentage of the total number of samples. From there, the methods that must be overwritten from the base Controller API are getControllerName(), computeControl(), and calculateSampledStateTrajectories(). Like in the other custom classes, getControllerName() returns the name of the new controller, i.e. CEM. Next, computeControl() is the main method of the Controller class. It takes the new initial state and calculates the new optimal control sequence from that starting position; an example is shown in Lst. 13. The basic steps are to move the shifted optimal control sequence and initial state to the GPU, generate control samples, run the samples through the Dynamics and Cost function, calculate the weights of each sample, use these weights to update the parameters of the sampling distribution, get the new optimal control sequence, and calculate the corresponding optimal state trajectory. In Lines 68, 69, 70, 71, 72, 73 and 74, we add an additional method, calcualteEliteSet(), to find the elite set and zero out the weights of every other sample specifically for CEM. The final method to overwrite, calculateSampledStateTrajectories(), is a method used to return a subset of sampled trajectories from the latest optimization round from the GPU to the CPU. Users can set a percentage of the sampled trajectories they would like returned with setPercentageSampledControlTrajectoriesHelper() and this method will generate state trajectories for those samples so that they can then be used for visualization in programs such as RViz. For this simple example, we have no visualization system to plug this data into so we leave this method empty for our CEM implementation.